For each function, determine the zeros. State the multiplicity of any multiple zeros. 2x^2-12x-32
Answer by ewatrrr(24785)

 
Hi,
2x^2-12x-32 = 0 x^2-6x-16 = 0 |factor
|The factor theorem states that a polynomial f(x) has a factor (x − k) if and only if f(k) = 0.
(x-8)(x+2) = 0
(x-8) = 0, x = 8
0r (x+2) = 0, x = -2